OpenStudy (anonymous):

Three letters are chosen at random from the word HEART and arranged in a row. Find the probability that: a) the letter H is chosen b) both vowels are chosen

Parth (parthkohli):

The probability of H being one of the three letters chosen = \(\Large {3 \over 5}\). The probability of H being selected in one of the three = \(\Large {1 \over 3}\) So, the answer to (a) will be \({\large{3 \over 5}} \times {\large {1 \over 3}}\).

Parth (parthkohli):

The probability that a vowel is chosen on the first pick is \(\Large {2 \over 5}\), and in the second pick = \(\Large {1 \over 4}\).

Parth (parthkohli):

So, the answer to (b) is \(\Large {2 \over 5} \times {1 \over 4}\)

OpenStudy (anonymous):

wait how is the probability of H being one of the three letters chosen=3/5? isn't it 1/5?

Parth (parthkohli):

You are right. Oopsie.

OpenStudy (anonymous):

nope, I am not actually, I am just trying to understand how you got it.

OpenStudy (anonymous):

and for part b the answer is 0.3 not 0.1 which you got.

Parth (parthkohli):

The probabilities of two independent events are multiplied.

Parth (parthkohli):

What? I can't be trusted any more. Sorry!

Parth (parthkohli):

\(5C3\)

OpenStudy (anonymous):

Now H is chosen so fix H first.. And the other two words can be select from 5 letters in: \[\large ^1C_1 \times ^5C_2 = ??\]

OpenStudy (anonymous):

I think it is also 10..

OpenStudy (anonymous):

My mind is not working I think..

OpenStudy (anonymous):

lol then I must not have one at all.

OpenStudy (anonymous):

See you have chosen 3 and put them in a row.. So they can be one out of the 5.. So parth is right to A part I guess..

OpenStudy (anonymous):

yes. he did get the correct answer which is 0.2 but so "The probability of three letters chosen from the 5 letter word Heart = 3/5. The probability of H being selected in one of the three = 1/3 and we just times them together 3/5 x 1/3 = 3/15 = 1/5 = 0.2

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Now from 5 we can choose 2 vowels in how many ways ???

OpenStudy (anonymous):

1/5 x 1/5 = 1/25?

OpenStudy (anonymous):

So: from 3 we have to choose 2 vowels now yes no??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So 1/9 I guess..

OpenStudy (anonymous):

@Ganpat help here in B part..

OpenStudy (ganpat):

probability for choosing 3 letters from 5 = 5C3 such that, both vowels r chosen = 2C2... as there r only two vowels in equation and remaining letter would be 3C1.. So the probability = ( 3C1 * 2C2 ) / 5C3 .. does this makes sense ??

OpenStudy (ganpat):

So the probability = ( 3C1 * 2C2 ) / 5C3 = ( 3 * 1) / (5 * 4 * 3)/(3 * 2) = 3/10 3/10, is probability for part b... recheck my calculations..

OpenStudy (cwrw238):

i think one way to do first part is the possibilities are HNN NHN NNH where N is not a H probability is 1/5 for first 4/5 * 1/4 for second and 4/5*3/4 * 1/3 for third add these up to get answer

OpenStudy (cwrw238):

= 1/5 + 4/20 + 12/60 = 3/5 - i'm not absolutely sure thats valid

OpenStudy (cwrw238):

does that make sense to you jays? probability is not my strong point.

OpenStudy (anonymous):

hmm not really and just a heads up that 3/5 is not the answer for part a.

OpenStudy (anonymous):

I guess I understand @Ganpat's method even though it's a bit confusing and I might forget how to apply how he did it to other similar questions.

OpenStudy (cwrw238):

do you have the answers?

OpenStudy (cwrw238):

correct answers?

OpenStudy (anonymous):

actually ur right, my bad.

OpenStudy (anonymous):

part a) 0.6 part b) 0.3

OpenStudy (cwrw238):

ok then ganpat was right about part b

OpenStudy (phi):

Three letters are chosen at random from the word HEART and arranged in a row. Find the probability that: a) the letter H is chosen We assume no replacement. Order does not matter (so combinations) there are 5C3 combinations in total. to find the number with an H, we can write down H _ _ where we have fixed the first choice to H. there are 4C2 ways to fill the 2 empty slots. Probability is \[\frac{\left(\begin{matrix} 4\\ 2\end{matrix}\right)}{\left(\begin{matrix}5 \\ 3\end{matrix}\right)}\] or 0.6

OpenStudy (phi):

Three letters are chosen at random from the word HEART and arranged in a row. Find the probability that: b) both vowels are chosen As above, the total number of patterns is 5C3= 10 if both vowels are chosen, we have A E _ (order does not matter) we can fill the last slot 3C1 different ways (5-2 leaves 3 letters to choose from) we get 3/10 = 0.3

OpenStudy (phi):

If we want to be pedantic, there is 1C1 way to choose A, 1C1 way to choose E, and 3C1 to pick the last letter, so the numerator is 1C1 * 1C1 * 3C1